Circular No. 8116

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF JUPITER AND SATURN Further to IAUC 8089, S. S. Sheppard reports the discovery of three

new satellites of Jupiter (S/2003 J 13, 14, and 15) and one new satellite of Saturn (S/2003 S 1), and B. Gladman (University of British Columbia) reports the discovery of three new satellites of Jupiter (S/2003 J 16, 17, and 18) — all from images obtained at Mauna Kea. Preliminary orbital elements for each object by B. G. Marsden appear on the MPEC references cited, as follows: S/2003 J 13, e = 0.41, i = 141^{o}, P = 738 days, 2003-G09; S/2003 J 14, e = 0.22, i = 141^{o}, P = 806 days, 2003-G10; S/2003 J 15, e = 0.11, i = 141^{o}, P = 667 days, 2003-G17; S/2003 J 16, e = 0.27, i = 149^{o}, P = 597 days, 2003-G18; S/2003 J 17, e = 0.19, i = 164^{o}, P = 690 days, 2003-G19; S/2003 J 18, e = 0.12, i = 146^{o}, P = 608 days, 2003-G20; S/2003 S 1, e = 0.33, i = 136^{o}, P = 989 days, 2003-G39. The Jupiter satellites have absolute magnitudes in the range H = 16.2–16.8; S/2003 S1 has H = 14.6.

COMET C/2003 G2 (LINEAR) L. Manguso, Lincoln Laboratory, Massachusetts Institute of Technol-

ogy, reports the LINEAR discovery of a comet with a 13′′ coma visible on Apr. 9–10 (discovery observation below). Following posting on the NEO Confirmation Page, the object was also reported to have cometary appearance by G. Hug (Eskridge, KS, 0.3-m reflector; diffuse with m_{1} = 16.6 on Apr. 9.4 UT and m_{1} = 17.3 on Apr. 10.4) and by A. C. Gilmore and P. M. Kilmartin (Mt. John University Observatory, 0.6-m reflector; diffuse on Apr. 11.6).

2003 UT  α_{2000}  δ_{2000}  m_{2}
Apr. 8.37712  14^{h}33^{m}50.^{s}98  −2^{o}22′32′′.0  18.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-G62.

T = 2003 Apr. 29.927 TT  ω = 190.^{o}680
Ω = 24.568 }2000.0
q = 1.55370 AU  i = 96.202
2003 April 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green